# O. Ragniscoa, F.Zullob

### v1

@@ 3.2 The case of an even Kernel
$\begin{split}&\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})=\sum_{n=1}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})-\sum_{n=1}\eta_{n}\exp(-Q\eta_{n}+\kappa_{n})=\\ &=2\sum_{n=1}\eta_{n}\exp(\kappa_{n})\sinh(Q\eta_{n})=\sum_{n\in\mathbb{Z}}\eta_{n}\exp(\kappa_{n})\sinh(Q\eta_{n})\end{split}$ (3.35)
@@
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(\kappa_{n})\sinh(Q\eta_{n}).$ (3.36)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\exp(\kappa_{n})\left(\cosh(Q\eta_{n})-1\right).$ (3.37)
Notice that we subtracted the constant term 1-1 to each term of the sequence in the Hamiltonian in order to ensure the convergence.
@@ 3.3 The case of an odd Kernel
$\begin{split}&\epsilon_{n}=\frac{\tilde{\epsilon}_{n}}{\sqrt{2N}},\quad|\tilde{\epsilon}_{n}|^{2}=1\quad\tilde{\epsilon}_{-n}=\tilde{\epsilon}_{n},\\ &\eta_{n}=\eta_{-n},\\ &\kappa_{n}=-\kappa_{-n}\end{split}$ (3.38)
@@
$\sum_{n=-N}^{N}\tilde{\epsilon}_{n}=o(\sqrt{N}),\quad\textrm{i.e.}\quad\lim_{N\to\infty}\sum_{n=-N}^{N}\epsilon_{n}=0.$ (3.39)
@@
$M(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}\exp(P\epsilon_{n})f(xs_{n}),$ (3.40)
where $f(x)$ is a continuous function of $x$ and $\epsilon_{n}$ is given in (3.38) and satisfies (3.39). Then, for any finite values of $x$ one has
$\lim_{N\to\infty}M(P,x,N)=Pf(0).$ (3.41)
Proof: we take the Taylor series of the exponential in (3.40) one has
$M(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}\left(1+P\epsilon_{n}+\frac{P^{2}}{2}\epsilon_{n}^{2}+\ldots\right)f(xs_{n}),$ (3.42)
@@
$P\sum_{n=-N}^{N}\epsilon_{n}^{2}f(xs_{n})=\frac{1}{2N}\sum_{n=-N}^{N}f(xs_{n}).$ (3.43)
From the Cauchy’s limit theorem on the arithmetic mean of sequences we see that this term converges to $Pf(0)$. It follows that all the powers of $\epsilon$ greater than $2$ does not give a contribution in the Taylor series expansion of $\exp(P\epsilon_{n})$ in (3.40). It remains only the contribution from the term linear in $\epsilon_{n}$. One has
$\sum_{n=-N}^{N}\epsilon_{n}f(xs_{n})=\frac{1}{\sqrt{2N}}\sum_{n=-N}^{N}\tilde{\epsilon}_{n}f(xs_{n}).$ (3.44)
Due to the condition (3.39), since $f(xs_{n})$ is bounded, the sum on the right hand side is $o(\frac{1}{\sqrt{N}})$ and it follows that the contribution from the term linear in $\epsilon$ in (3.40) vanishes $\square$.
@@
$\dot{Q}=\sum_{n=-N}^{N}\epsilon_{n}\exp(P\epsilon_{n}+Q\eta_{n}+\kappa_{n})$ (3.45)
@@
$\dot{Q}=P.$ (3.46)
Let us now look at the equation of motion for $P$. We notice that Proposition (3.3) is still valid if $\epsilon_{n}$ is as given in (3.20), the proof being the same. So we can directly write
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}).$ (3.47)
@@
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n})\cosh(\kappa_{n}).$ (3.48)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\exp(Q\eta_{n})\left(\cosh(\kappa_{n})-1\right).$ (3.49)
At this point it is also possible to take a flow that is a combination of the Hamiltonian containing the even Kernel (3.37), say $\mathcal{H}_{e}$, and the Hamiltonian containing the odd Kernel, say $\mathcal{H}_{o}$:
$\alpha\mathcal{H}_{e}+(1-\alpha)\mathcal{H}_{o}=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\left(\alpha\exp(Q\eta^{o}+\kappa^{e}_{n})+(1-\alpha)\exp(Q\eta^{e}+\kappa^{o}_{n})-1\right).$ (3.50)
where the superscripts on the sequences remind if they are odd or even. Let us now give a Remark about the equilibrium positions.
Remark 3.6
@@
$\dot{Q}=P,\quad\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}),$ (3.51)
where the sequences $\epsilon_{n}$ and $\eta_{n}$ have opposite parity with respect to $\kappa_{n}$. The above system has always just one equilibrium position, say $(Q,P)=(Q_{0},0)$ if the sequence $\eta_{n}$ has not a definite sign. The motion is bounded and the equilibrium position is a center.
[…]
From another point of view we can say that the system (3.51) is Hamiltonian with the Hamiltonian given by
$\mathcal{H}=T+V=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\left(\exp(Q\eta_{n}+\kappa_{n})-1\right).$ (3.52)
Figure 1: Plot of the curves defined by $\dot{Q}=0$ (red) and $\dot{P}=0$ (black) dividing the plane $(Q,P)$ in four regions, each having a definite sign for $\dot{Q}$ and $\dot{P}$.
Remark 3.7
If the sequence $\eta_{n}$ has just one sign, then the system (3.51) has one equilibrium position if $\eta_{n}\geq 0$ and has no equilibrium position (open orbits) if $\eta_{n}\leq 0$.
The previous Remark follows by looking at the properties of the system (3.51) when $\eta_{n}\geq 0$ or $\eta_{n}\leq 0$, as did for the Remark (3.6)
Remark 3.8
The equations of motion (3.51) holds true also if the sequences $\eta_{n}$ and $\kappa_{n}$ belong to $\ell^{1}$. The Propositions (3.1), (3.3) and (3.4) indeed hold also for $\eta_{n}$ and $\kappa_{n}$ belonging to $\ell^{1}$.
An example of sequences in $\ell^{1}$ will be given in the subsection (5.2)
@@ 4.1 General outlook
$\frac{\partial y}{\partial t}=\epsilon(x)-\int_{-\infty}^{\infty}dx^{\prime}(\epsilon(x)\eta(x^{\prime})-\eta(x)\epsilon(x^{\prime}))\exp[y(x^{\prime})]$ (4.2)
@@
Even in this case, for the existenceexistence of the equations of motion one needs that the argument of the exponential be uniformly bounded on the whole real line, while $\epsilon(x),\eta(x)$ must be absolutely integrable, and then square integrable, on the whole real line. Of course these requirements are "a fortiori" satisfied by $S^{\infty}$ functions.
@@
$\dot{P}=1-\int_{-\infty}^{\infty}dx\eta(x)\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]$ (4.4)
$\dot{Q}=\int_{-\infty}^{\infty}dx\epsilon(x)\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]$ (4.5)
Equations (4.4,4.5) can be written in canonical form, namely
$\dot{P}=-\frac{\partial H}{\partial Q}$ (4.6)
$\dot{Q}=\frac{\partial H}{\partial P}$ (4.7)
@@
$H=-Q+\int_{-\infty}^{\infty}dx\exp(P\epsilon(x)+Q\eta(x)+\kappa)$ (4.8)
As in the infinite discrete case, the Hamiltonian (4.8) does not seem to exist, even if $\kappa(x),~\epsilon(x),~\eta(x)$ are uniformly bounded and rapidly decreasing functions. However, we remark that, analogously to the infinite discrete case, if the exponential is uniformly bounded on the whole real line, then the improper integral can be made convergent by just subtracting to the integrand its asymptotic limit, i.e. $1$.
@@
AA is a skew-symmetric Hilbert-Schmidt operator since its kernel is square summable if so are the functions $\epsilon(x)$ and $\eta(x)$); its kernel is in fact the kernel of the Poisson Bracket between two functionals $F[y],G[y]$:
@@
We can apply to $iA$ all the theorems holding for compact self-adjoint operators [11] on a Hilbert space. In particular, it has no residual spectrum, its eigenvectors, including those associated with its kernel, form a complete (orthonormal) set in $L^{2}({\mathbb{R}})$, its discrete spectrum is either finite or given by a decreasing sequence accumulating to 0, and its norm is a point of the spectrum. Indeed, we have more, because the kernel $A(x,x^{\prime})$ is a separable kernel of rank 2, and consequently the operator $A$ has only two nonzero complex conjugated eigenvalues.
It is remarkable, though probably obvious, that the results for eigenvalues and eigenfunctions holding in the discrete case migrate with no essential modifications to the continuum case.
@@
$\psi^{\pm}(x,\omega)=(1/\lambda_{\pm})(\eta(x)\alpha_{\pm}-\epsilon(x)\beta_{\pm}))$ (4.9)
The nonzero eigenvalues are given by:
$\lambda_{\pm}\doteq\pm i\omega=\pm i{\sqrt{|\epsilon|^{2}|\eta|^{2}-(\epsilon,\eta)^{2}}}$ (4.10)
@@
$|\eta|^{2}\alpha-(\eta,\epsilon)\beta=\lambda\beta$ (4.11)
$(\epsilon,\eta)\alpha-|\epsilon|^{2}\beta=\lambda\alpha$ (4.12)
Once solved for $\alpha,\beta$ the system (4.11), (4.12), in terms of variables $\epsilon,\eta$ we get for the eigenfunctions $\psi_{\pm}(x.\omega)\doteq u(x,\omega)\pm iv(x,\omega)$ the following expressions:
$u(x,\omega)=\epsilon(x)/|\epsilon|$ (4.13)
$w(x,\omega)=(\epsilon(x)(\epsilon,\eta)-\eta(x)|\epsilon|^{2})/(\omega|\epsilon|)$ (4.14)
As in the finite $N$ case, $u(x,\omega)$ and $w(x,\omega)$ have norm one and are mutually orthogonal. Accordingly., the kernel $A(x,x^{\prime})$ of the integral operator $A$ can be written in the form:
$A(x,x^{\prime})=i\omega(u(x,\omega)w(x^{\prime},\omega)-u(x^{\prime},\omega)w(x,\omega))$ (4.15)
Now, let us assume that the system enjoys an equilibrium configuration, namely that there exist at least one point $P^{0},Q^{0}$ in the $P,Q$-plane such that the function $z(x)\doteq\exp[P^{0}\epsilon(x)+Q^{0}\eta(x)+\kappa(x)]$ fulfills the properties $(z,\epsilon)=0$, $(z,\eta)=1$, and ask whether there exist small oscillations around that equilibrium configuration. To this aim, we set $P=P^{0}+\delta_{1},\quad Q=Q^{0}+\delta_{2}$, and consider a first order expansion of the exponential, setting
$\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]\simeq z(x)(1+\delta_{1}\epsilon(x)+\delta_{2}\eta(x))$
Accordingly, we get the following linear system of differential equations for the unknowns $\delta_{1},\delta_{2}$:
$\dot{\delta}_{1}=-{\cal A}\delta_{1}-{\cal B}\delta_{2},$ (4.16)
$\dot{\delta}_{2}={\cal C}\delta_{1}+{\cal A}\delta_{2}.$ (4.17)
@@
${\cal A}=\int_{-\infty}^{\infty}dxz(x)\epsilon(x)\eta(x)$ (4.18)
${\cal B}=\int_{-\infty}^{\infty}dxz(x)\eta(x)^{2}$ (4.19)
${\cal C}=\int_{-\infty}^{\infty}dxz(x)\epsilon(x)^{2}$ (4.20)
@@
The nature of the system (4.16-4.17) is fixed by the solution of the secular equation:
$\lambda^{2}-{\cal A}^{2}+{\cal B}{\cal C}=0$ (4.21)
@@
${\cal A}^{2}+{\cal B}{\cal C}=(1/2)\int_{-\infty}^{\infty}dx\int_{-\infty}^{\infty}dx^{\prime}z(x)z(x^{\prime})[\epsilon(x)\eta(x^{\prime})-\eta(x)\epsilon(x^{\prime})]^{2}$ (4.22)
where (4.22) is the trace of the square of the matrix $M$
$M=\begin{pmatrix}-{\cal A}&-{\cal}B\\ {\cal C}&{\cal A}\end{pmatrix}$ (4.23)
hence, the linearized motion will be periodic around the centrum $P^{0},Q^{0}$ with period
$\omega={\sqrt{-{\cal A}^{2}+{\cal B}{\cal C}}}$

### v2

@@ 3.2 The case of an even Kernel
$\begin{split}&\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})=\sum_{n=1}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})-\sum_{n=1}\eta_{n}\exp(-Q\eta_{n}+\kappa_{n})=\\ &=2\sum_{n=1}\eta_{n}\exp(\kappa_{n})\sinh(Q\eta_{n})=\sum_{n\in\mathbb{Z}}\eta_{n}\exp(\kappa_{n})\sinh(Q\eta_{n})\end{split}$ (47)
@@
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(\kappa_{n})\sinh(Q\eta_{n}).$ (48)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\exp(\kappa_{n})\left(\cosh(Q\eta_{n})-1\right).$ (49)
Notice that we subtracted the constant term 1-1 to each term of the sequence in the Hamiltonian in order to ensure the convergence. The same situation occurs in the infinite Toda lattice [20, 19] and in the integrable infinite Volterra lattice obtained from the infinite Toda by the reduction $B_{n}=0$ [9, 20]. This is the analogous to the renormalization procedure in field theory when the vacuum state energy is subtracted to the Hamiltonian.
@@ 3.3 The case of an odd Kernel
$\begin{split}&\epsilon_{n}=\frac{\tilde{\epsilon}_{n}}{\sqrt{2N}},\quad|\tilde{\epsilon}_{n}|^{2}=1\quad\tilde{\epsilon}_{-n}=\tilde{\epsilon}_{n},\\ &\eta_{n}=\eta_{-n},\\ &\kappa_{n}=-\kappa_{-n}\end{split}$ (50)
@@
$\sum_{n=-N}^{N}\tilde{\epsilon}_{n}=o(\sqrt{N}),\quad\textrm{i.e.}\quad\lim_{N\to\infty}\sum_{n=-N}^{N}\epsilon_{n}=0.$ (51)
@@
$M(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}\exp(P\epsilon_{n})f(xs_{n}),$ (52)
where $f(x)$ is a continuous function of $x$ and $\epsilon_{n}$ is given in (50) and satisfies (51). Then, for any finite values of $x$ one has
$\lim_{N\to\infty}M(P,x,N)=Pf(0).$ (53)
Proof. Let us take the Taylor series of the exponential in (52):
$M(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}\left(1+P\epsilon_{n}+\frac{P^{2}}{2}\epsilon_{n}^{2}+\ldots\right)f(xs_{n}),$ (54)
@@
$P\sum_{n=-N}^{N}\epsilon_{n}^{2}f(xs_{n})=\frac{1}{2N}\sum_{n=-N}^{N}f(xs_{n}).$ (55)
From the Cauchy’s limit theorem on the arithmetic mean of sequences we see that this term converges to $Pf(0)$. It follows that all the powers of $\epsilon$ greater than $2$ do not give a contribution in the Taylor series expansion of $\exp(P\epsilon_{n})$ in (52). It remains only the contribution from the term linear in $\epsilon_{n}$. One has
$\sum_{n=-N}^{N}\epsilon_{n}f(xs_{n})=\frac{1}{\sqrt{2N}}\sum_{n=-N}^{N}\tilde{\epsilon}_{n}f(xs_{n}).$ (56)
Due to the condition (51), since $f(xs_{n})$ is bounded, the sum on the right hand side is $o(\frac{1}{\sqrt{N}})$ and it follows that the contribution from the term linear in $\epsilon$ in (52) vanishes $\square$.
@@
$\dot{Q}=\sum_{n=-N}^{N}\epsilon_{n}\exp(P\epsilon_{n}+Q\eta_{n}+\kappa_{n})$ (57)
@@
$\dot{Q}=P.$ (58)
Let us now look at the equation of motion for $P$. We notice that Proposition (3.3) is still valid if $\epsilon_{n}$ is as given in (32), the proof being the same. So we can directly write
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}).$ (59)
@@
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n})\cosh(\kappa_{n}).$ (60)
@@
$\mathcal{H}=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\exp(Q\eta_{n})\left(\cosh(\kappa_{n})-1\right).$ (61)
Let us now give a Remark about the equilibrium positions.
Remark 3.6.
@@
$\dot{Q}=P,\quad\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}),$ (62)
where the sequences $\epsilon_{n}$ and $\eta_{n}$ have opposite parity with respect to $\kappa_{n}$. The above system has always just one equilibrium position, say $(Q,P)=(Q_{0},0)$ if the sequence $\eta_{n}$ has not a definite sign. In this case the motion is bounded and the equilibrium position is a center.
Indeed the function defining $\dot{P}$ in (62) is a decreasing function of $Q$ as it follows immediately by taking its derivative with respect to $Q$. Further, for $Q\to$-$\infty$, $\dot{P}\to+\infty$ (we assume that there will be always at least one $\eta_{n}$ negative), whereas for $Q\to+\infty$, $\dot{P}\to$-$\infty$ (we assume that there will be always at least one $\eta_{n}$ positive). It follows that the equation $\dot{P}=0$ has just one zero. Further, the motion must be bounded and periodic. Notice that the equilibrium point is stable but not asymptotically stable: this easily follows from the linearization of the system (62) around the equilibrium point, that is a center. The fact that the motion is bounded and periodic follows from a global analysis of the orbits in the plane $(Q,P)$, or, more easily, by looking at the potential in the Hamiltonian defining the Hamiltonian system (62), given by
$\mathcal{H}=T+V=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\left(\exp(Q\eta_{n}+\kappa_{n})-1\right).$ (63)
The potential
$V=-Q+\sum_{n\in\mathbb{Z}}\left(\exp(Q\eta_{n}+\kappa_{n})-1\right)$ (64)
has one global minimum $Q=Q_{0}$. Indeed, by assuming that the sequence $\eta_{n}$ has not a definite sign, $V$ approaches $\infty$ for $Q\to\pm\infty$. Also, its derivative has a unique zero, as shown by Remark (3.6). This zero is a minimum for the potential $V$, as can be easily seen by taking the second derivative of $V$. So $V(Q)$ is an infinite potential well and the motion is bounded and periodic. In Figure (1) we report the nullclines corresponding to $\dot{Q}=0$ and $\dot{P}=0$ in red and the arrows displaying the direction of the flow in the four region of the plane $(Q,P)$ defined by the nullclines.
Figure 1: Plot of the curves defined by $\dot{Q}=0$ and $\dot{P}=0$ (red) dividing the plane $(Q,P)$ in four regions, each having a definite sign for $\dot{Q}$ and $\dot{P}$.
Remark 3.7.
If the sequence $\eta_{n}$ has just one sign, then the system (62) has one equilibrium position if $\eta_{n}\geq 0$ and has no equilibrium position (open orbits) if $\eta_{n}\leq 0$.
The previous Remark follows by looking at the properties of the system (62) when $\eta_{n}\geq 0$ or $\eta_{n}\leq 0$, as did for the Remark (3.6). Indeed if $\eta_{n}\leq 0$ for all $n$ the equation defining $\dot{P}$, i.e.
$\dot{P}=1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}),$ (65)
is always positive and $P$ increases monotonically in time, resulting in a unbounded motion. Vice versa, if $\eta_{n}\geq 0$ then the equation $\dot{P}=0$ has just one solution since $1-\sum_{n\in\mathbb{Z}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n})$ is a decreasing is a decreasing function of $Q$ and goes from $1$ for $Q\to$-$\infty$ to -$\infty$ for $Q\to+\infty$. The equilibrium position is again a center in this case and the orbits are bounded and periodic.
Remark 3.8.
The equations of motion (62) holds true also if the sequences $\eta_{n}$ and $\kappa_{n}$ belong to $\ell^{1}$. The Propositions (3.1), (3.3) and (3.4) indeed hold also for $\eta_{n}$ and $\kappa_{n}$ belonging to $\ell^{1}$.
An example of sequences in $\ell^{1}$ will be given in the subsection (5.2). To end this Section, we give a Remark about the possibility to get a mixed odd/even Hamiltonian.
Remark 3.9.
It is possible to take a flow that is a combination of the Hamiltonian containing the even Kernel (49), say $\mathcal{H}_{e}$, and the Hamiltonian containing the odd Kernel, say $\mathcal{H}_{o}$:
$\alpha\mathcal{H}_{e}+(1-\alpha)\mathcal{H}_{o}=\frac{P^{2}}{2}-Q+\sum_{n\in\mathbb{Z}}\left(\alpha\exp(Q\eta_{n}^{o}+\kappa^{e}_{n})+(1-\alpha)\exp(Q\eta_{n}^{e}+\kappa^{o}_{n})-1\right).$ (66)
where the superscripts on the sequences remind if they are odd or even. If the sequences $\eta_{n}^{o}$ and $\eta_{n}^{e}$ have not a definite sign, the potential
$V(Q)=-Q+\sum_{n\in\mathbb{Z}}\left(\alpha\exp(Q\eta_{n}^{o}+\kappa^{e}_{n})+(1-\alpha)\exp(Q\eta_{n}^{e}+\kappa^{o}_{n})-1\right)$ (67)
has just one minima for $\alpha\in[0,1]$ and is an infinite potential well, so that the corresponding motion is bounded and periodic. The proof is the same as after Remark (3.6) and is omitted.
@@ 4.1 General outlook
$\frac{\partial y}{\partial t}=\epsilon(x)-\int_{-\infty}^{\infty}dx^{\prime}(\epsilon(x)\eta(x^{\prime})-\eta(x)\epsilon(x^{\prime}))\exp[y(x^{\prime})]$ (69)
@@
Even in this case, for the existenceexistence of the equations of motion one needs that the argument of the exponential be uniformly bounded on the whole real line, while $\epsilon(x),\eta(x)$ must be absolutely integrable, and then square integrable, on the whole real line. Of course these requirements are ”a fortiori” satisfied by $S^{\infty}$ functions.
@@
$\dot{P}=1-\int_{-\infty}^{\infty}dx\eta(x)\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]$ (71)
$\dot{Q}=\int_{-\infty}^{\infty}dx\epsilon(x)\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]$ (72)
Equations (71,72) can be written in canonical form, namely
$\dot{P}=-\frac{\partial H}{\partial Q}$ (73)
$\dot{Q}=\frac{\partial H}{\partial P}$ (74)
@@
$H=-Q+\int_{-\infty}^{\infty}dx\exp(P\epsilon(x)+Q\eta(x)+\kappa(x))$ (75)
As in the infinite discrete case, the Hamiltonian (75) does not seem to exist, even if $\kappa(x),~\epsilon(x),~\eta(x)$ are uniformly bounded and rapidly decreasing functions. However, we remark that, analogously to the infinite discrete case, if the exponential is uniformly bounded on the whole real line, then the improper integral can be made convergent by just subtracting to the integrand its asymptotic limit, i.e. $1$.
@@
AA is a skew-symmetric Hilbert-Schmidt operator since its kernel is square summable if so are the functions $\epsilon(x)$ and $\eta(x)$; its kernel is in fact the kernel of the Poisson Bracket between two functionals $F[y],G[y]$:
@@
We can apply to $iA$ all the theorems holding for compact self-adjoint operators [10] on a Hilbert space. In particular, it has no residual spectrum, its eigenvectors, including those associated with its kernel, form a complete (orthonormal) set in $L^{2}({\mathbb{R}})$, its discrete spectrum is either finite or given by a decreasing sequence accumulating to 0, and its norm is a point of the spectrum. Indeed, we have more, because the kernel $A(x,x^{\prime})$ is a separable kernel of rank 2, and consequently the operator $A$ has only two non-zero complex conjugated eigenvalues.
It is remarkable that the results for eigenvalues and eigenfunctions holding in the discrete case migrate with no essential modifications to the continuum case.
@@
$\psi^{\pm}(x,\omega)=(1/\lambda_{\pm})(\eta(x)\alpha_{\pm}-\epsilon(x)\beta_{\pm}))$ (76)
The non-zero eigenvalues are given by:
$\lambda_{\pm}\doteq\pm i\omega=\pm i{\sqrt{|\epsilon|^{2}|\eta|^{2}-(\epsilon,\eta)^{2}}}$ (77)
@@
$|\eta|^{2}\alpha-(\eta,\epsilon)\beta=\lambda\beta$ (78)
$(\epsilon,\eta)\alpha-|\epsilon|^{2}\beta=\lambda\alpha$ (79)
Once solved for $\alpha,\beta$ the system (78), (79), in terms of variables $\epsilon,\eta$ we get for the eigenfunctions $\psi_{\pm}(x.\omega)\doteq u(x,\omega)\pm iv(x,\omega)$ the following expressions:
$u(x,\omega)=\epsilon(x)/|\epsilon|$ (80)
$w(x,\omega)=(\epsilon(x)(\epsilon,\eta)-\eta(x)|\epsilon|^{2})/(\omega|\epsilon|)$ (81)
As in the finite $N$ case, $u(x,\omega)$ and $w(x,\omega)$ have norm one and are mutually orthogonal. Accordingly, the kernel $A(x,x^{\prime})$ of the integral operator $A$ can be written in the form:
$A(x,x^{\prime})=i\omega(u(x,\omega)w(x^{\prime},\omega)-u(x^{\prime},\omega)w(x,\omega))$ (82)
4.1.1 Linearization around the eventual equilibrium configuration.
Now, let us assume that the system enjoys an equilibrium configuration, namely that there exist at least one point $P^{0},Q^{0}$ in the $P,Q$-plane such that the function $z(x)\doteq\exp[P^{0}\epsilon(x)+Q^{0}\eta(x)+\kappa(x)]$ fulfils the properties $(z,\epsilon)=0$, $(z,\eta)=1$, and ask whether there exist small oscillations around that equilibrium configuration. To this aim, we set $P=P^{0}+\delta P,\quad Q=Q^{0}+\delta Q$, and consider a first order expansion of the exponential, setting
$\exp[P\epsilon(x)+Q\eta(x)+\kappa(x)]\simeq z(x)(1+\delta P\epsilon(x)+\delta Q\eta(x))$
Accordingly, we get the following linear system of differential equations for the unknowns $\delta P,\delta Q$:
$\dot{\delta P}=-{\cal A}\delta P-{\cal B}\delta Q,$ (83)
$\dot{\delta Q}={\cal C}\delta P+{\cal A}\delta Q.$ (84)
@@
${\cal A}=\int_{-\infty}^{\infty}dxz(x)\epsilon(x)\eta(x)$ (85)
${\cal B}=\int_{-\infty}^{\infty}dxz(x)\eta(x)^{2}$ (86)
${\cal C}=\int_{-\infty}^{\infty}dxz(x)\epsilon(x)^{2}$ (87)
@@
The nature of the system (83-84) is fixed by the solution of the secular equation:
$\lambda^{2}-{\cal A}^{2}+{\cal B}{\cal C}=0$ (88)
@@
${\cal A}^{2}+{\cal B}{\cal C}=(1/2)\int_{-\infty}^{\infty}dx\int_{-\infty}^{\infty}dx^{\prime}z(x)z(x^{\prime})[\epsilon(x)\eta(x^{\prime})-\eta(x)\epsilon(x^{\prime})]^{2}$ (89)
where (89) is the trace of the square of the matrix $M$
$M=\begin{pmatrix}-{\cal A}&-{\cal}B\\ {\cal C}&{\cal A}\end{pmatrix}$ (90)
hence, the linearized motion will be periodic around the centrum $(P^{0},Q^{0})$ with period
$\omega={\sqrt{-{\cal A}^{2}+{\cal B}{\cal C}}}.$